# Diagnosing and Mitigating Semantic Inconsistencies in Wikidata’s Classification Hierarchy

### v1

@@ 3.3.3 Semantic Drift Computation
Step 1: Mean Parent Embedding
@@
$\vec{p}_{\text{mean}}=\frac{1}{n}\sum_{i=1}^{n}\vec{p}_{i}$ (1)
Step 2: Raw Semantic Drift
@@
$\text{drift}_{\text{raw}}=1-\cos(\vec{e},\vec{p}_{\text{mean}})$ (2)
Step 3: Drift Adjustment
@@
$\text{drift}_{\text{adj}}=\text{drift}_{\text{raw}}\cdot\log(n+1)$ (3)
@@
Step 4: Thresholding
Entities with $\text{drift}_{\text{adj}}\geq 0.60$ are flagged as semantically inconsistent. The selection of this threshold is empirically justified by the global distribution of drift scores across different structural complexities, as detailed in our Results section (see Figure [3]).
Fig. 3: Heatmap of entity counts distributed by adjusted drift score and parent group
Figure [3]: Heatmap of entity counts distributed by adjusted drift score and parent group. The distribution patterns of semantic drift vary significantly based on an entity’s structural complexity, as defined by its number of parent classes.
The heatmap illustrates that the vast majority of entities, those with two or fewer parents, are heavily concentrated in low-drift bins (adjusted drift ¡ 0.6). For this group, the entity count drops sharply beyond the [0.4, 0.6) bin, indicating that scores above this range are statistically uncommon and likely represent genuine semantic inconsistencies. Conversely, entities with a higher number of parents (3-6 and  6) exhibit distributions shifted towards higher drift scores, with their counts peaking in the [0.6, 1.0) range.
The value 0.6 serves as a critical inflection point; it effectively identifies outliers for the largest group of structurally simple entities while remaining sensitive to the inherently higher drift values of more complex entities. Therefore, this threshold provides a reasonable and data-driven basis for flagging potential semantic drift across the knowledge graph.
3.3.4 Aggregation and Domain Analysis
We recursively assign each entity to its pseudo-root class (a node that is never a child of any other node) by performing BFS propagation from root nodes over $\mathcal{E}_{\text{clean}}$. This mapping allows us to compute drift statistics per root class.
For each root class $r$, we compute:
  • Total number of assigned entities ($\text{cnt}_{r}$)
  • Average drift ($\text{avg\_drift}_{r}$)
  • 90th percentile drift ($\text{p90}_{r}$)
  • High-drift ratio: % of entities with $\text{drift}_{\text{adj}}\geq 0.60$
We visualize distributional drift per segment (violin plots) and per domain (bar plots for top-20 root classes), enabling both qualitative inspection and quantitative benchmarking of structural coherence across knowledge domains.
4 Results
4.1 CME Detection Part
In the first stage, we selected different domains to test the method of using graphical judgment to determine the mixed use of P31/P279 relationships. The final results are as follows: we found that the severity of the mixed-use issue varies across different fields, and the results of our random tests are consistent with previous research findings, which proves that our method in this phase is effective.
Fig. 4: Stage one Result
In order to first gain an understanding of Wikidata as a whole, we have compiled some data: There are 4.92 million entities in Wikidata that are used as classes (linked to other entities through the P279 relationship). To avoid semantic drift, we set conditions for filtering: entities with a P279 count equal to 1 and a P31 count less than or equal to 2 (i.e., excluding entities that have at most one redundant connection after abstract definitions). After filtering, there are still 3.5 million entities remaining. If we exclude entities that do not contain instances, there are ultimately 43,065 entities left as meaningful pure classes. By analyzing the tree structure of pure classes, we finally obtain 24,793 entities that can be split into separate entity trees. These entity trees contribute about 10 million instances, accounting for 10% of the Wikidata data
Fig. 5: clean classes with their percentage of instances
Table 1: Representative Examples of Semantic Drift Scores
@@ 4.1 CME Detection Part
To provide a more concrete understanding of what these drift scores represent, Table [1] presents a qualitative analysis of representative entities sampled from low, medium, and high-drift regions of the distribution.
4.2 Semantic Drift Results
Fig. 6: Semantic Drift in different group of entities
@@ 4.2 Semantic Drift Results
5 Interface
@@
Fig. 7: interface for entity check system
6 Discussion
6.1 Redefining Semantic Inconsistency: From ”Error” to ”Risk”
A primary contribution of this research is the reframing of semantic inconsistency not as a binary error to be corrected, but as a quantifiable risk to be managed. Traditional quality assessments often apply a dichotomous correct/incorrect label, a paradigm that proves insufficient for a complex, community-curated knowledge graph like Wikidata. As evidenced by Dadalto et al. (2022) [2], many taxonomic anti-patterns arise from the inherently multifaceted nature of real-world concepts, suggesting that simplistic, one-sided corrections are often inappropriate.
Our ”semantic risk” framework offers a more nuanced alternative. It acknowledges that inconsistencies exist but quantifies them as a spectrum of risk. This approach respects Wikidata’s core principle of ”plurality of facts” [8] while providing actionable insights for downstream applications. For example, an entity with a high-risk score might be flagged for careful handling in high-precision tasks like commonsense reasoning, yet its connections could remain valuable in others. This perspective also informs the debate on redundant edges. Rather than wholesale removal, our risk model suggests a filtering mechanism: high-risk redundant links (e.g., those connecting to semantically distant parents) should be prioritized for review, while low-risk ones (e.g., hierarchical shortcuts) could be preserved to serve as useful heuristics, as suggested by Yan et al. (2020) [10].
6.2 The Interplay and Tension Between Structure and Semantics
@@
6.3 Generality and Limitations of the Methodology
@@
6.4 Practical Implications for Wikidata and Downstream Applications
@@ 8 Future Works
8.1 Expanded Validation Datasets
@@
8.2 Benchmark Comparison Analysis
@@
8.3 Multi-dimensional Evaluation Metrics

### v2

@@ 6.3. Semantic Drift Computation
Mean Parent Embedding
@@
(1) $\vec{p}_{\text{mean}}=\frac{1}{n}\sum_{i=1}^{n}\vec{p}_{i}$
Raw Semantic Drift
@@
(2) $\text{drift}_{\text{raw}}=1-\cos(\vec{e},\vec{p}_{\text{mean}})$
Drift Adjustment
@@
(3) $\text{drift}_{\text{adj}}=\text{drift}_{\text{raw}}\cdot\log(n+1)$
@@
Thresholding
Entities with $\text{drift}_{\text{adj}}\geq 0.60$ are flagged as semantically inconsistent. The selection of this threshold is empirically justified by the global distribution of drift scores across different structural complexities, as detailed in Figure6.
Figure 6. Heatmap of Entity Counts Distributed by Adjusted Drift Score and Parent Group
Figure6 shows that the distribution of semantic drift varies substantially with an entity’s structural complexity, measured by the number of its parent classes.
Entities with two or fewer parents, which constitute the vast majority of the dataset, are highly concentrated in low-drift regions (adjusted drift ¡ 0.6). For this group, the entity count declines sharply beyond the [0.4, 0.6) bin, indicating that higher drift values are statistically rare and are more likely to correspond to genuine semantic inconsistencies. In contrast, entities with more complex structures (3–6 parents and more than 6 parents) exhibit distributions that are clearly shifted toward higher drift scores, with their frequencies peaking in the [0.6, 1.0) range.
The value 0.6 therefore emerges as a critical inflection point. It effectively isolates outliers among structurally simple entities while remaining sensitive to the inherently higher drift levels associated with structurally complex entities. As such, this threshold provides a reasonable and data-driven criterion for flagging potential semantic drift across the knowledge g
6.4. Case Study and Results
To provide a more concrete understanding of what these drift scores represent, Table2 presents a qualitative analysis of representative entities sampled from low, medium, and high-drift regions of the distribution.
Table 2. Representative Examples of Semantic Drift Scores
@@ 6.4. Case Study and Results
Figure 7. Semantic Drift in Different Group of Entities
Figure 7 presents a Kernel Density Estimation (KDE) plot that reveals a distinct correlation between structural complexity and semantic drift. The distribution of drift scores exhibits a clear monotonic shift to the right as the number of parent classes increases:
  • High Coherence in Simple Structures (Group A): The blue curve, representing entities with $\leq 2$ parents, is characterized by a sharp, narrow distribution concentrated in the low-drift region (peak $\approx 0.35$). The density drops precipitously after 0.60, confirming that structurally simple entities typically maintain high semantic cohesion. This sharp drop-off empirically validates our selection of 0.60 as a robust threshold for detecting anomalies in the majority of Wikidata entities.
  • Inherent Divergence in Complex Structures (Group E): Conversely, the green curve (entities with >6>6 parents) shows a significantly broader and flatter distribution with a peak around 1.5. This ”long-tail” behavior indicates that entities with heavy distinct inheritance chains are inherently more prone to semantic divergence. The high drift scores here reflect both the cumulative semantic distance of diverse parent concepts and the logarithmic penalty term in our metric.
  • Transitional Behavior (Group C): The orange curve serves as a bridge, showing that moderate structural complexity (363-6 parents) leads to a predictable increase in semantic ambiguity, centered around the 0.6-0.9 range.
7. Interface-Supported
@@
Figure 8. Interface for Entity Check System
8. Discussion
8.1. Redefining Semantic Inconsistency: From ”Error” to ”Risk”
A primary contribution of this research is the reframing of semantic inconsistency not as a binary error to be corrected, but as a quantifiable risk to be managed. Traditional quality assessments often apply a dichotomous correct/incorrect label, a paradigm that proves insufficient for a complex, community-curated knowledge graph like Wikidata. As evidenced by Dadalto et al. (2022) (Dadalto2024EvidenceOL), many taxonomic anti-patterns arise from the inherently multifaceted nature of real-world concepts, suggesting that simplistic, one-sided corrections are often inappropriate.
Our ”semantic risk” framework offers a more nuanced alternative. It acknowledges that inconsistencies exist but quantifies them as a spectrum of risk. This approach respects Wikidata’s core principle of ”plurality of facts” (Wikidata) while providing actionable insights for downstream applications. For example, an entity with a high-risk score might be flagged for careful handling in high-precision tasks like commonsense reasoning, yet its connections could remain valuable in others. This perspective also informs the debate on redundant edges. Rather than wholesale removal, our risk model suggests a filtering mechanism: high-risk redundant links (e.g., those connecting to semantically distant parents) should be prioritized for review, while low-risk ones (e.g., hierarchical shortcuts) could be preserved to serve as useful heuristics, as suggested by Yan et al. (2020) (Yan2020LearningCK).
8.2. The Interplay and Tension Between Structure and Semantics
@@
8.3. Generality and Limitations of the Methodology
@@
8.4. Practical Implications for Wikidata and Downstream Applications
@@ 10. Future Works
10.1. Expanded Validation Datasets
@@
10.2. Benchmark Comparison Analysis
@@
10.3. Multi-dimensional Evaluation Metrics